OpenStudy (anonymous):

Which equation best represents the problem? You have traveled 200 miles in two and a half hours. How fast were you driving?

OpenStudy (lilg132):

80 mph

OpenStudy (anonymous):

200/2.5 miles/hr

OpenStudy (anonymous):

D = (200)(2.5) 200 = R (2.5) 2.5 = R (200) 2.5 = 200 T

OpenStudy (anonymous):

2.5 * speed = 200

OpenStudy (lilg132):

M / T = x mph

OpenStudy (lilg132):

Miles divided by Time = x Mph 200 divided by 2.5 = 80 mph Please dont forget to give me a medal or good answer :D

OpenStudy (anonymous):

So it is B